Consider the following economy of exchange. There are two goods and two consumers. The two goods are called tillip and quillip and the two consumers are called 1 and 2.

Consumer 1 has the utility function U1(t, q) = . 4ln(t)+ . 6ln(q) (where t is the amount of tillip and q is the amount of quillip).

Consumer 2 has the utility function U2(t,q)=. 5ln(t)+. 5ln(q).

Consumer 1 is equipped with 10units of quillip and tillip each. Consumer 2 is equipped with 10 quillip units and 5 tillip units.

a) What is the Walrasian equilibrium of this economy? ( If there is more than one equilibrium write them all down).
Answer #1

Ui(tiq)- 0-4 ln t +0.6 ln q entot + engub Since stility ultiq = 0.4 0.6 - en t q is an ordinal concept t t et qui oo Similary2 2 2 2 2 =P 29= 3t p q= ЗЕР Substituting in Income constraint Pt+lq < 10 p +10 3pt 10C1+p) pt tapt 5 pt = 10(1+p) 2 t = to CU₂ (t,q)= to 5 Consumer 2 0.5 max to 5quos Sub.to pt tlq = 1045p - MUE Р Muq aulot subg P 오 I 7) t q=pt Substitute in incomeTotal demand of quilep- Total endowment of quelp 10+52 + 6(1+p) = 20 2 5+2.5p +6+ 6p = 20 11 +8.5p = 20 8.5 p=9 p=9 8-5 P. 16